# Ambient Noise Full Waveform Inversion with Neural Operators

### v2

@@ 5 Discussion
In this study, we have demonstrated the functionality of neural operators in realistic full waveform ambient noise inversion through a 2D case. Provided sufficient computational resources, the proposed method is expected to work in 3D without fundamental changes, while providing increased computational advantages over conventional methods. We aim to provide some plug-and-play models pre-trained to high accuracy for the community to apply to other regions on Earth. Because our model is trained within the framework of supervised learning, the fundamental assumption is that training and testing data follow the same distribution. Therefore, the model predictions are expected to fail where the velocity parameters deviate significantly from the training distribution. We show the training distributions of $V_{P}$ and $V_{S}$ associated with depth in Figure 13, allowing users to decide whether to directly use our model or perform additional training. Even in cases where re-training is necessary, a pre-trained model enables transfer learning that can greatly improve learning efficiency (Yosinski et al., 2014). Another aspect is that the model is not expected to extrapolate to larger physical domains than which it was trained on ($80$ km length by $40$ km depth) or to frequencies outside the preset band of interest.
Figure 13: Training distributions of $V_{P}$ and $V_{S}$ associated with depth.
@@
In performing FWI, we use gradient descent with AD instead of training a reverse model that directly maps from observed waveforms to velocity parameters. Although such end-to-end approaches are even faster (Moseley et al., 2020; Wu & Lin, 2019), they provide only a single solution without supervision on the data misfit. Such approaches are arguably not ideal for inverse problems because they are ill-posed and non-unique, unless they are combined with uncertainty quantification methods. We define the misfit function for inversion as the MSE in the frequency domain for convenient use with HNO, but it would be feasible and worthwhile to explore different misfit functions in either the frequency or time domain (Bozdağ et al., 2011; Métivier et al., 2018; Fournier & Oldenburg, 2019; Sambridge et al., 2022) to better tackle the non-uniqueness and cycle-skipping problems. However, the AD method requires a differentiable loss function, which is not compatible with, e.g., the cross-correlation time shift commonly used in FWI, unless some smooth approximation is applied.
@@
The rapid developments in compute capabilities and machine learning research have positioned data-driven approaches as a promising new perspective on FWI. On one hand, data-driven approaches can greatly accelerate modeling and inversion while suffering less from typical gradient-based approximations in physics-driven approaches. On the other hand, they rely on data provided by physics-driven approaches and, therefore, will also benefit from any advancements in those approaches. In this study, we integrate these two types of approaches by first generating training data with a physics-driven solver and then training a data-driven model as a surrogate solver. Another way is to incorporate physics into the training loss, making it possible to remove the external solver (Li et al., 2024; Raissi et al., 2019; Rasht-Behesht et al., 2022; Ren et al., 2024; Song & Alkhalifah, 2021). While training with physics loss alone is challenging, combining it with a data loss has the potential to improve generalization ability and further benefit inversion.
@@ 6 Conclusions
We present the first application of neural operators in full waveform ambient noise inversion using several real seismic datasets from linear nodal arrays distributed across the northern LA basins. We show that a Helmholtz Neural Operator trained on random velocity fields can generalize to realistic velocity structures, while demonstrating robustness to random noise. In performing FWI with automatic differentiation, the neural operator is two orders of magnitude faster than the conventional SEM and does not require a separate adjoint source calculation. In comparison to the adjoint-state method relying on the Born approximation, our method has not linearized any gradients and is therefore less prone to cycle-skipping problems. The tomography results from real data align with multiple previous studies and can be obtained in a plug-and-play manner. We provide a trained model that can be directly applied to different regions, as long as the velocity parameters fall within the training distribution.
@@
The node data can be accessed from the IRIS Data Management Center (BASIN, 2018), and the code, models, and processed data are available at https://github.com/caifeng-zou/ANFWI_HNO (Zou, 2025).
@@ Acknowledgements
ZER acknowledges financial support from the David and Lucile Packard Foundation.
@@
Figure S1: Loss curves. The batch size is $256$. We use an Adam optimizer with a learning rate of $0.001$ and a scheduler that decays the learning rate by half every $30$ epochs.
Figure S2: FWI results for noise-free data using the adjoint-state method with the same 1D initialization as in Figure 5.
Figure S3: Inverted $V_{S}$ for SB1 with minimum inter-station distances of $5$ km (Figure 8a) and $10$ km.
@@
Table S1: Details in each FNO layer.
@@
1 32 (lifting dimension) 64 (192, 96) (96, 48) GELU
2 64 128 (128, 64) (64, 32) GELU
3 128 256 (64, 32) (32, 16) GELU
4 256 512 (32, 16) (16, 8) GELU
5 512 256 (64, 32) (16, 8) GELU
6 512 128 (128, 64) (32, 16) GELU
7 256 64 (192, 96) (64, 32) GELU
8 128 64 (256, 128) (96, 48) GELU

### v3

@@ 4 Discussion
In this study, we have demonstrated the functionality of neural operators in realistic ambient noise full waveform inversion through a 2D case. Provided sufficient computational resources, the proposed method is expected to work in 3D without fundamental changes, while providing increased computational advantages over conventional methods. Because our model is trained within the framework of supervised learning, the fundamental assumption is that training and testing data follow the same distribution. While this assumption seldom holds strictly in practice, it serves as a reference for identifying when model predictions are likely to fail — for example, in cases where the velocity parameters deviate significantly from the training distribution. We show the training distributions of $V_{P}$ and $V_{S}$ as a function of depth in Figure 12, allowing users to decide whether to directly use our model or perform additional training. Even in cases where re-training is necessary, a pre-trained model enables transfer learning that can greatly improve learning efficiency (Yosinski et al., 2014). Another aspect is that the model is not expected to extrapolate to larger physical domains than which it was trained on (808080 km length by 404040 km depth) or to frequencies outside the preset band of interest.
Figure 12: Training distributions of $V_{P}$ and $V_{S}$ as a function of depth.
@@
In performing FWI, we use gradient descent with AD instead of training a reverse model that directly maps from observed waveforms to velocity parameters. Although such end-to-end approaches are even faster (Kazei et al., 2021; Moseley et al., 2020; Wu & Lin, 2019), they provide only a single solution without supervision on the data misfit. Such approaches are arguably not ideal for inverse problems because they are ill-posed and non-unique, unless they are combined with uncertainty quantification methods. We define the misfit function for inversion as the MSE in the frequency domain for convenient use with HNO, but it would be feasible and worthwhile to explore different misfit functions in either the frequency or time domain (Bozdağ et al., 2011; Métivier et al., 2018; Fournier & Oldenburg, 2019; Sambridge et al., 2022; Sun & Alkhalifah, 2019) to better tackle the non-uniqueness and cycle-skipping problems. It is important to note, however, that the AD method requires a differentiable loss function, or some soft approximation to address discontinuity (Liu et al., 2023b).
@@
The rapid developments in compute capabilities and machine learning research have positioned data-driven approaches as a promising new perspective for FWI. On one hand, data-driven approaches can greatly accelerate modeling and inversion. On the other hand, they rely on data provided by physics-driven approaches and, therefore, will also benefit from any advancements in those approaches. In this study, we integrate these two types of approaches by first generating training data with a physics-driven solver and then training a data-driven model as a surrogate solver. Another way is to incorporate physics into the training loss, making it possible to remove the external solver (Li et al., 2024; Raissi et al., 2019; Rasht-Behesht et al., 2022; Ren et al., 2024; Song & Alkhalifah, 2021). While training with physics loss alone is challenging, combining it with a data loss has the potential to improve generalization ability and further benefit inversion.
@@ 5 Conclusions
We present the first application of neural operators in ambient noise full waveform inversion using several real seismic datasets from linear nodal arrays distributed across the northern LA basins. We show that a Helmholtz Neural Operator trained on random velocity fields can generalize to realistic velocity structures, while demonstrating robustness to random noise. In performing FWI with automatic differentiation, the neural operator is two orders of magnitude faster than the conventional spectral element method and eliminates the need for manual gradient derivation. The HNO-AD method, built on PyTorch, offers enhanced flexibility to escape poor local minima. The tomography results from real data align with multiple previous studies and can be obtained with minimal additional effort. We provide a trained model that can be directly applied to different regions, as long as the velocity parameters fall within the training distribution.
@@
The node data can be accessed from the IRIS Data Management Center (BASIN, 2018). The network codes are 4M and 6J. The code, models, and processed data are available at https://github.com/caifeng-zou/ANFWI_HNO (Zou, 2025).
@@ Acknowledgements
ZER acknowledges financial support from the David and Lucile Packard Foundation. CZ and RWC acknowledge support from NSF EAR-2105358 and EAR-2438773. FCL acknowledges financial support from the Statewide California Earthquake Center (SCEC) based on Award Number DE-SC0016520 from the U.S. Department of Energy and NSF EAR-2438772.
@@
Figure S1: Loss curves. The batch size is 256256256. We use an Adam optimizer with a learning rate of 0.0010.0010.001 and a scheduler that decays the learning rate by half every 303030 epochs.
Figure S2: FWI results for noise-free data using the HNO-AD and adjoint-state methods with the same initialization. The HNO-AD method uses mini-batch Adam (not yet supported by SALVUS), while the adjoint-state method uses full-batch L-BFGS. The result from HNO-AD is noisier than that in Figure 6, because all frequencies are handled together here.
@@
Table S1: Details in each neural operator layer.
@@
FNO 1 32 (lifting dimension) 64 (192, 96) (96, 48) GELU
FNO 2 64 128 (128, 64) (64, 32) GELU
FNO 3 128 256 (64, 32) (32, 16) GELU
FNO 4 256 512 (32, 16) (16, 8) GELU
FNO 5 512 256 (64, 32) (16, 8) GELU
FNO 6 512 128 (128, 64) (32, 16) GELU
FNO 7 256 64 (192, 96) (64, 32) GELU
FNO 8 128 64 (256, 128) (96, 48) GELU
Layer Input channel Output channel Dimension Hidden neurons Activation
GNO 100 96 (256, 1) (96, 96) GELU